# Comments on the Emergence of 4D Topological Amplitudes in M-Theory

### 19 Mar 2026 v1

@@ Abstract
The M-theoretic Emergence Proposal claims that all of the terms in the low-energy action arise from quantum effects. After reviewing the current status of this proposal, we focus on four-dimensional compactifications with $N=2$ supersymmetry, where kinetic terms are encoded in topological string amplitudes, such as the prepotential $\mathcal{F}_{0}$. Evidence for the emergence of such terms was provided recently, where in particular it was shown that the classical cubic term in $\mathcal{F}_{0}$ can be obtained by integrating out the light towers of states in the M-theory limit, using a novel regularization of the infinite sum over Gopakumar/Vafa invariants. We address two issues that were left open. First, we show that the regularization can be equivalently performed in complex structure moduli space and in Kähler moduli space. Second, we extend the proposed regularization to the linear terms in the one-loop prepotential $\mathcal{F}_{1}$.
@@ 2.1 Preliminaries
$\mathcal{F}_{1}(t)={2\pi i}\left(\frac{c_{2}}{24}t+\frac{1}{2\pi i}\sum_{\beta\in H_{2}(X,\mathbb{Z})}\left(\frac{1}{12}\alpha_{0}^{\beta}+\alpha_{1}^{\beta}\right){\rm Li}_{1}(e^{2\pi i\beta\cdot t})\right)\,$, (3)
@@
$\mathcal{F}_{1}(t)=\sum_{(\beta,n)\neq(0,0)}\left(\frac{1}{12}\alpha_{0}^{\beta}+\alpha_{1}^{\beta}\right)\int_{0}^{\infty}\frac{ds}{s}e^{sZ_{\beta,n}}\,.$ (5)
@@
$\begin{split}\frac{1}{2}\sum_{n=1}^{\infty}\alpha_{n}n^{3}=-\sum_{n,k=1}^{\infty}\alpha_{n}n^{3}=-\lim_{\tau\to 0}\sum_{n,k=1}^{\infty}\alpha_{n}n^{3}e^{-2\pi nk\tau}=-\lim_{\tau\to 0}\sum_{n=1}^{\infty}\alpha_{n}n^{3}\frac{q^{n}}{1-q^{n}}\end{split}$ (7)
@@
$\begin{split}\frac{1}{2}\sum_{n=1}^{\infty}\alpha_{n}n^{3}=-\lim_{\tau\to 0}\sum_{n}^{\infty}\alpha_{n}n^{3}\frac{q^{n}}{1-q^{n}}=-\lim_{t\to 0}\left[(\partial_{t}^{3}\mathcal{F}_{0}|_{\rm weak}-\kappa_{ttt})\right]\end{split}$ (8)
@@ 3.2 Limit in Kähler moduli space
as $\Delta t$ becomes small. Thus, adding the subleading terms necessary to cancel the $\log|\log u|$ contribution and match the $1/\log u$ piece does not alter the leading divergent behavior that had already been fixed. If one wanted to find the expansion up to $1/\log^{2}u$-order, the same strategy would apply. Higher orders in all the expansions would need to be added, together with the subleading terms $\log|\log\Delta t|/(\Delta t\log^{4}\Delta t)$ and $1/(\Delta t\log^{4}\Delta t)$ which would not affect the previous orders.
@@ 4 Regularization scheme for the 1-loop prepotential ${\cal F}_{1}$
$\mathcal{F}_{1}=-\sum_{(\beta,n)\neq(0,0)}\left(\frac{1}{12}\alpha_{0}^{\beta}+\alpha_{1}^{\beta}\right)\left(\gamma_{E}+\log(-\epsilon Z_{n})+\mathcal{O}(\epsilon)\right)\,$, (30)
@@
[…]
@@
$C_{t}^{(0)}=-(2\pi i)\lim_{t\to t_{c}}\left[\left(\left.\partial_{t}\mathcal{F}_{1}\right|_{\rm weak}-\frac{c_{2}}{24}\right)-{\rm Div}+\frac{1}{24\pi i}\partial_{t}\log\epsilon(t)\right]\,.$ (34)
@@ 4.1 Limit in complex structure moduli space
$\mathcal{F}_{1}=\frac{1}{2}\log\left[e^{\left(3+h_{1,1}-\frac{\chi}{12}\right)K}\det(G^{-1})|f|^{2}\right]\,$, (36)
where $f$ is a holomorphic function, the holomorphic anomaly, which should be determined by imposing the appropriate behavior at the singular points of the moduli space. Focusing now on the mirror quintic, the LCS behavior $\mathcal{F}_{1}\rightarrow\frac{50}{24}(t+\bar{t})$, together with imposing regularity at the orbifold point $\psi=0$, fixes $f$ to be
@@
$\partial_{t}\mathcal{F}_{1}=\frac{\partial\psi}{\partial t}\partial_{\psi}\log\left[e^{\frac{31}{3}K}\det\left(G_{\psi\bar{\psi}}^{-1}\right)^{\frac{1}{2}}\psi^{\frac{62}{6}}(1-\psi^{-5})^{-\frac{1}{12}}\right]\,.$ (38)
@@
$\left.\partial_{t}\mathcal{F}_{1}\right|_{\rm weak}=\frac{1}{u\log u}\left(\sum_{n,k=0}^{\infty}\frac{(u\log u)^{n}}{\log^{k}u}\tilde{a}_{(n,k)}(u)\right)\,$, (39)
@@ 4.2 Limit in Kähler moduli space
$\partial_{t}\mathcal{F}_{1}\bigr|_{\rm weak}(u)=\frac{1}{u\log u}\left(\tilde{a}_{(0,0,0)}+\tilde{a}_{(0,1,0)}\frac{1}{\log u}+\mathcal{O}\big(\log^{-2}u\big)\right)\,$, (44)
@@
$\partial_{t}\mathcal{F}_{1}(\Delta t)=\alpha_{1}\frac{1}{\Delta t}+\alpha_{2}\frac{1}{\Delta t\log\Delta t}+\mathcal{O}\left(\frac{1}{\Delta t\log^{2}\Delta t}\right)$ (48)

### 31 Mar 2026 v2

@@ Abstract
The M-theoretic Emergence Proposal claims that all of the terms in the low-energy action arise from quantum effects. After reviewing the current status of this proposal, we focus on four-dimensional compactifications with $N=2$ supersymmetry, where kinetic terms are encoded in topological string amplitudes, such as the prepotential $\mathcal{F}_{0}$. Evidence for the emergence of such terms was provided recently, where in particular it was shown that the classical cubic term in ${\cal F}_{0}$ can be obtained by integrating out the light towers of states in the M-theory limit, using a novel regularization of the infinite sum over Gopakumar/Vafa invariants. We address two issues that were left open. First, we show that the regularization can be equivalently performed in complex structure moduli space and in Kähler moduli space. Second, we extend the proposed regularization to the linear terms in the one-loop prepotential $\mathcal{F}_{1}$.
@@ 2.1 Preliminaries
$\mathcal{F}_{1}(t)={2\pi i}\ \left(\frac{c_{2}}{24}t-\frac{1}{2\pi i}\sum_{\beta\in H_{2}(X,\mathbb{Z})}\left(\frac{1}{12}\alpha_{0}^{\beta}+\alpha_{1}^{\beta}\right){\rm Li}_{1}(e^{2\pi i\beta\cdot t})\right)\,$, (3)
@@
$\mathcal{F}_{1}(t)=-\sum_{(\beta,n)\neq(0,0)}\left(\frac{1}{12}\alpha_{0}^{\beta}+\alpha_{1}^{\beta}\right)\int_{0}^{\infty}\frac{ds}{s}e^{sZ_{\beta,n}}\,.$ (5)
@@
$\begin{split}\frac{1}{2}\sum_{n=1}^{\infty}\alpha_{0}^{n}n^{3}=-\sum_{n,k=1}^{\infty}\alpha_{0}^{n}n^{3}=-\lim_{\tau\to 0}\sum_{n,k=1}^{\infty}\alpha_{0}^{n}n^{3}e^{-2\pi nk\tau}=-\lim_{\tau\to 0}\sum_{n=1}^{\infty}\alpha_{0}^{n}n^{3}\frac{q^{n}}{1-q^{n}}\end{split}$ (7)
@@
$\begin{split}\frac{1}{2}\sum_{n=1}^{\infty}\alpha_{0}^{n}n^{3}=-\lim_{\tau\to 0}\sum_{n}^{\infty}\alpha_{0}^{n}n^{3}\frac{q^{n}}{1-q^{n}}=-\lim_{t\to 0}\left[(\partial_{t}^{3}\mathcal{F}_{0}|_{\rm weak}-\kappa_{ttt})\right]\end{split}$ (8)
@@ 3.2 Limit in Kähler moduli space
as $\Delta t$ becomes small. Thus, adding the subleading terms necessary to cancel the $\log|\log u|$ contribution and match the $1/\log u$ piece does not alter the leading divergent behavior that had already been fixed. If one wanted to find the expansion of (19) up to order $1/(u\log^{5}u)$, the same strategy would apply. Higher orders in all the expansions would need to be added, together with the subleading terms $\log|\log\Delta t|/(\Delta t\log^{4}\Delta t)$ and $1/(\Delta t\log^{4}\Delta t)$ which would not affect the previous orders.
@@ 4 Regularization scheme for the 1-loop prepotential ${\cal F}_{1}$
$\mathcal{F}_{1}=\sum_{(\beta,n)\neq(0,0)}\left(\frac{1}{12}\alpha_{0}^{\beta}+\alpha_{1}^{\beta}\right)\left(\gamma_{E}+\log(-\epsilon Z_{\beta,n})+\mathcal{O}(\epsilon)\right)\,$, (30)
@@
$\begin{split}\mathcal{F}_{1}^{D0,D2}&=\sum_{\beta\neq 0}\left(\frac{1}{12}\alpha_{0}^{\beta}+\alpha_{1}^{\beta}\right)\left(-\pi i\,\beta t+\log\big(1-e^{2\pi i\beta t}\big)\right)\,,\\[2.84544pt] \mathcal{F}_{1}^{D0}&=\frac{1}{12}\log\mu\,.\end{split}$ (31)
@@
$C_{t}^{(0)}=-(2\pi i)\lim_{t\to t_{c}}\left[\left(\left.\frac{1}{2\pi i}\partial_{t}\mathcal{F}_{1}\right|_{\rm weak}-\frac{c_{2}}{24}\right)-{\rm Div}+\frac{1}{24\pi i}\partial_{t}\log\epsilon(t)\right]\,.$ (34)
@@ 4.1 Limit in complex structure moduli space
$\mathcal{F}_{1}=\pi i\log\left[e^{\left(3+h_{1,1}-\frac{\chi}{12}\right)K}\det(G^{-1})|f|^{2}\right]\,$, (36)
where $f$ is a holomorphic function, the holomorphic anomaly, which should be determined by imposing the appropriate behavior at the singular points of the moduli space. Focusing now on the mirror quintic, the LCS behavior $\mathcal{F}_{1}\rightarrow 2\pi i\ \frac{50}{24}(t+\bar{t})$, together with imposing regularity at the orbifold point $\psi=0$, fixes $f$ to be
@@
$\partial_{t}\mathcal{F}_{1}=2\pi i\ \frac{\partial\psi}{\partial t}\partial_{\psi}\log\left[e^{\frac{31}{3}K}\det\left(G_{\psi\bar{\psi}}^{-1}\right)^{\frac{1}{2}}\psi^{\frac{62}{6}}(1-\psi^{-5})^{-\frac{1}{12}}\right]\,.$ (38)
@@
$\left.\partial_{t}\mathcal{F}_{1}\right|_{\rm weak}=\frac{2\pi i}{u\log u}\left(\sum_{n,k=0}^{\infty}\frac{(u\log u)^{n}}{\log^{k}u}\tilde{a}_{(n,k)}(u)\right)\,$, (39)
@@ 4.2 Limit in Kähler moduli space
$\partial_{t}\mathcal{F}_{1}\bigr|_{\rm weak}(u)=\frac{2\pi i}{u\log u}\left(\tilde{a}_{(0,0,0)}+\tilde{a}_{(0,1,0)}\frac{1}{\log u}+\mathcal{O}\big(\log^{-2}u\big)\right)\,$, (44)
@@
$\partial_{t}\mathcal{F}_{1}(\Delta t)=\alpha_{1}\frac{2\pi i}{\Delta t}+\alpha_{2}\frac{2\pi i}{\Delta t\log\Delta t}+\mathcal{O}\left(\frac{1}{\Delta t\log^{2}\Delta t}\right)$ (48)